OFIS Notes from 16 Sep, 2010

Past CS tasks

  • Daniel: Research Book work: Talking with MWG and PaF


  • Marlon will move data updates from dev to prod (GSO and Office of Research).
  • Vu has nearly finished code updates for that data; needs to do report upgrades.
  • Daniel will report next week on faculty concerns. 24 faculty have saved their profiles (out of 82).
  • Marlon: Office of Research is switching systems; unknown when will be in production.
  • Beth working on better data-processing via Eng Dean's office; discussions with Alan George about how we're doing all sorts of things to get around HR limitations; Alan G will be meeting with Janet Passmore, head of HR; Beth meeting with a direct report of hers, responsible for change management.

Future Tasks

  • Daniel: research book stuff, ct'd.; report for next week about faculty concerns; test Vu's contributions import system

-- DanielAllen - 16 Sep 2010

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2010-09-16 - DanielAllen
This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2024 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback